a sphere of radius R=40 and mass M=7.2kg. calculate the moment of inertia about the axis of rotation at a distance of R from the centre of sphere.


1
Expert's answer
2021-07-14T10:16:37-0400

"I=\frac 23 mr^2+mr^2=\frac 53 mr^2,"

"I=\frac 53\cdot 7.2\cdot 0.4^2=1.92~kg\cdot m^2."